OpenStudy (anonymous):

Which expressions are completely factored? Choose all answers that are correct. Question 4 options: 16y^5 + 12y^3 = 4y^3(4y^2 + 3) 18y^3 – 6y = 3y(6y^2 – 2) 20y^7 + 10y^2 = 5y(4y^6 + 2y) 32y^10 – 24 = 8(4y^10 – 3)

OpenStudy (babyx3boo):

Is it me or are all of them correct? o.o

OpenStudy (anonymous):

lol i was thinking that earlier when i saw so i left and felt awkward @babyx3boo

OpenStudy (babyx3boo):

lol then that must mean we are correct xD @Migoz300

OpenStudy (anonymous):

probably does then lol

OpenStudy (babyx3boo):

@NiaLiveQueenB our guess is that each one of them is right

OpenStudy (anonymous):

Which statements are true? Choose exactly two answers that are correct. Question 3 options: 4x^2 + 10x = 2x(2x + 5) 8x^3 – 6x = 2x^3(4 – 3x^3) 30x^4 – 12x^3 = 6x^3(5x – 2) 100x^3 + 5 = 5x^3(20x + 1)

OpenStudy (anonymous):

ok isn't this one a and c

OpenStudy (babyx3boo):

You are correct :)

OpenStudy (anonymous):

and btw ist a and d not all of it is right

OpenStudy (anonymous):

well it was an assumption

OpenStudy (anonymous):

yea thnks btw

OpenStudy (babyx3boo):

and also they all came the same to me when i did it lol but oops sorry :/

OpenStudy (anonymous):

same soooo my bad B